OpenStudy (venomblast):

How to find Secant lines if point P(6,-2) lies on a curve y=2/(5-x) If Q is the point (x,2/(5-x)) and find the slope of mPQ

OpenStudy (anonymous):

The slope of a function between two points is \[\frac{\Delta y}{\Delta x}=\frac{y_2-y_1}{x_2-x_1}~~?\] In this case, \((x_1,y_1)=(6,-2)\) and \((x_2,y_2)=\left(x,\dfrac{2}{5-x}\right)\) Plug these coordinates in and see if you can simplify it.

OpenStudy (venomblast):

Ohh o. the (x, 2/(5-x) ) confused me

OpenStudy (anonymous):

typo there \[\large \frac{\frac{2}{5-x}+2}{x-6}\]

OpenStudy (venomblast):

ok

OpenStudy (anonymous):

might be easiest algebra if you multiply top and bottom by \(5-x\) leave the denominator in factored form

OpenStudy (venomblast):

i thought it was \[d1-d2/t1-t2\]

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

\[y_1=\frac{2}{5-x}, y_2=-2, x_1=x, x_2=6\]

OpenStudy (venomblast):

so is it reverse?

OpenStudy (venomblast):

will i get this -2x+12/x-6

OpenStudy (venomblast):

im confused